OpenStudy (anonymous):

Solve the initial value equation for x(t): dx/dt = 1-x^2, x(0)=3

OpenStudy (anonymous):

I got \[t(x) = \frac{ 1 }{ 2 }\ln \frac{ |x+1| }{ |x-1| }+3\] After integrating.

OpenStudy (anonymous):

Book says the answer is: \[x(t) = \frac{ 2+e ^{-2t} }{ 2-e ^{-2t} }\] I don't know how they got there.

OpenStudy (anonymous):

How did you integrate?

OpenStudy (anonymous):

Apparently wrong, hold on while I redo everything from the integration down.

OpenStudy (anonymous):

Err, I got excited for nothing, integration was right, so my partial fractions were: \[dx = \frac{ 1 }{ -2(x-1) }+\frac{ 1 }{ 2(x-1) }\]

OpenStudy (anonymous):

why is your x-1 on top?

OpenStudy (anonymous):

Can you confirm the partial fractions above?

OpenStudy (anonymous):

I used dx / (1-x^2)= A / x+1 + b / ( 1-X)

OpenStudy (anonymous):

does that help ? B=1/2, A=1/2

OpenStudy (anonymous):

The solutions used (1-x) too, but I don't know why I can't use -(x-1)(x+1) and get the same answer.

OpenStudy (anonymous):

Every time I try to think on my own, wrong, wrong, wrong.

OpenStudy (anonymous):

you certainly can , but then you need to care for the " - ve sign " and use properties of log to rewrite : ) you may be correct, DONT ALWAYS COMPARE your answer with book. If you do need to compare, work backwords to your answer.

OpenStudy (anonymous):

Can you arrange it my way and try to work though it with me?

OpenStudy (anonymous):

I would say THERE IS NO POINT in writing that -ve sign "your way" , it may complicate matters. Save time : )

OpenStudy (anonymous):

But you understand its frustrating not being able to complete a problem in your own way right?

OpenStudy (anonymous):

I dont see reason for frustration, what matters is the approach to solve the problem

OpenStudy (anonymous):

did you get 1/2 * [ ln |(1+x)]| +1/2*[ln|(1-x)|]=t+ K

OpenStudy (anonymous):

- 1/2 ln(1-x)

OpenStudy (anonymous):

put x=3 and t=0 1/2 * [ln(1+3)]+1/2*[ln|(1-3)|]=0+K

OpenStudy (anonymous):

I'm going to upload a pic of my sheet, hold on a sec.

OpenStudy (anonymous):

Okie

OpenStudy (anonymous):

the work is all fine, until the very last step, you DO NOT need that negative sign, you have already taken care of it

OpenStudy (anonymous):

ok, so now I'll find c

OpenStudy (anonymous):

if you say -ln(a)+ln(b)= ln (b/a), then there is NO NEED to write it as - ln (b/a)

OpenStudy (anonymous):

Ya, I get it, I make a lot of dumb errors

OpenStudy (anonymous):

before you find C, make sure you correct your mistakes, a mistake of an incorrect sign will give incorrect values of C

OpenStudy (anonymous):

So it appears C is 1?

OpenStudy (anonymous):

and dont compare your answer with book, you are on right track, just find value of C and leave it as it is.

OpenStudy (anonymous):

I hope that helps, am taking off here

OpenStudy (anonymous):

Nooooooooo, I need you